Trend-Line Inc. has been growing at a rate of 6% per year and is expected to continue to do so indefinitely. The next dividenURGENT!!
Answer #1

a.

The formula to calculate price of stock is given below:

Price\, of\, stock\, =\, \frac{Expected\, dividend}{r\, -\, g}

Here,

r is rate of return.

g is growth rate.

Substitute $8 for expected dividend, 10% for r and 6% for g,

Price\, of\, stock\, = \frac{8}{0.10\, -\, 0.06} =\frac{8}{0.04} =200

Hence, current selling price is $200.

b.

If zero-plowback strategy is used by the firm, then it could pay the dividend in perpetuity.

The value of assets in place is calculated below:

Value\, of\, asset\,in\: place = \frac{Dividend}{Rate of return} =\frac{12}{0.10} =120

c.

The value which remains after deducting value of assets in place from value of stock must be value due to growth opportunity.

The part value due to growth opportunity is calculated below:

Part\, value\, due\, to\, growth\, opportunity = Value\, of\; stock\: -\: Value\, of\, assets\, in\, place =\, 200-120 =80Hence, part of share value due to growth opportunity is $80.